# Identification by non-Gaussianity in structural threshold and smooth transition vector autoregressive models

### v2

@@ B.6 Details on the empirical application
The ML estimation of the models, residual diagnostics, and estimation of the GIRFs are carried out with the accompanying R package sstvars (Virolainen, 2024) in which the introduced methods have been implemented. The R package sstvars also contains the dataset studied in the empirical application to facilitate reproduction of our results. In the ML estimation, we consider only solutions that can be deemed local maximums of the log-likelihood function.
In the below residual based model diagnostics, we have not labelled the residual series by any of the variables, because identification with respect to the ordering of the columns of $B_{2}$ is weak (see Section 3.5). One the other hand, changing the ordering of the columns of $B_{y,t}$ changes the ordering of the variables in the density function $t_{1}(I_{d,i}^{\prime}B_{y,t}^{-1}(y_{t}-\mu_{t});0,1,\nu_{i})$, thus, changing the ordering of the residuals in the given regime. The residual time series can, therefore, be labelled only for the variables whose structural shocks have been labelled. Since we labelled the CPU shock as the first shock in Section 5.3, the first residual series constitutes the residuals of CPUI. Since the other structural shocks were not labelled, the other residual series cannot be labelled for any of the variables.
@@
To study whether our model adequately captures the autocorrelation structure of the data, we depict the sample autocorrelation functions (ACF) and cross-correlation functions (CCF) of the residuals for the first $24$ lags in Figure B.1. As the figure shows, there is not much autocorrelation left in the residuals, but there is a particularly large correlation coefficient (CC) in the first lag of the CCFs between the fourth and fifth residual series. There are also some moderate sized CCFs in the ACFs of the third residual series, but these are not particularly large. Since rest of the CCs are reasonable, our LSTVAR model seems to capture the autocorrelation structure of the data reasonably well.
@@
In order to study the model’s adequacy to capture the conditional heteroskedasticity in the data, we depict the sample ACFs and CCFs of squared standardized residuals for the first $24$ lags in Figure B.2. The figure shows that there are several moderate and relative large CCs in the ACFs and CCFs of some of the residual series. Thus, there is some but not particularly severe inadequacies in terms of capturing the conditional heteroskedasticity.
@@
The series of standardized residuals, presented in the top panels of Figure B.3, also show some remaining heteroskedasticity and several outliers, many of which are accommodated by the heavy tailed $t$-distributions. Finally, the quantile-quantile-plots based on the independent $t$-distributions with mean zero, variance one, and degrees-of-freedom given by the degrees-of-freedom parameter estimate are presented in the bottom panel of Figure B.3. They show that the marginal distributions of the series are relatively well captured, although the distributions of two of the residual series are slightly asymmetric. Nonetheless, in our view, the overall adequacy of the model seems reasonable enough.
@@
Figure B.1: Auto- and crosscorrelation functions of the residuals of the fitted two-regime fifth-order LSTVAR model for the lags $0,1,...,24$. The lag zero autocorrelation coefficients are omitted, as they are one by convention. The dashed lines are the $95\%$ bounds ±$\pm 1.96/\sqrt{T}$ for autocorrelations of IID observations.
@@
Figure B.2: Auto- and crosscorrelation functions of the squared standardized residuals of the fitted two-regime fifth-order LSTVAR model for the lags $0,1,...,24$. The lag zero autocorrelation coefficients are omitted, as they are one by convention. The dashed lines are the $95\%$ bounds ±$\pm 1.96/\sqrt{T}$ for autocorrelations of IID observations.
@@
Figure B.3: Standardized residual time series and quantile-quantile-plots of the fitted two-regime fifth-order LSTVAR model. The quantile-quantile plots are based on $t$–distributions with zero mean, variance one, and degrees of freedom given by the ML estimate of the degrees of freedom parameter.

### v3

@@ Appendix E Details on the empirical application
The estimation of the models, residual diagnostics, and computation of the GIRFs are carried out with the accompanying R package sstvars (Virolainen, 2025a) in which the introduced methods have been implemented. The R package sstvars also contains the dataset studied in the empirical application. In the below residual based model diagnostics, we have not labelled the residual series by any of the variables, because labelling them effectively requires labelling the corresponding structural shocks, and only the CPU shock has been labelled.
@@
To study whether our model adequately captures the autocorrelation structure of the data, we depict the sample autocorrelation functions (ACF) and cross-correlation functions (CCF) of the (non-standardized) residuals for the first $24$ lags in Figure E.1. There is not much autocorrelation left in the residuals, but there is a particularly large correlation coefficient (CC) in the lag zero CCFs between the fourth and fifth residual series. In addition, there are a few moderately sized CCs in the ACFs and CCFs of some of the residual series, but the vast majority of the CCs are small. Hence, with the exception of the single large CC in the lag zero CCFs of the fourth and fifth residual series, our LSTVAR model appears to capture the autocorrelation structure of the data reasonably well.
@@
To study our model’s adequacy to capture the conditional heteroscedasticity in the data, we depict the sample ACFs and CCFs of squared standardized residuals for the first $24$ lags in Figure E.2. There is a very large CC in the lag zero CCFs between the second and third residual series, which, however, disappears if the COVID-19 residual from March 2020 is excluded (not shown). In addition, there are several sizeable CCs in the ACFs and CCFs of some of the residual series. Thus, there are some inadequacies in capturing the conditional heteroscedasticity, but with the exception of the large CC in CCFs between the second and third residual series (due to the extreme COVID-19 volatility), the inadequacies do not seem particularly severe.
@@
Finally, the quantile-quantile-plots based on the independent skewed $t$-distributions with mean zero, variance one, and degrees-of-freedom as well as skewness parameter values given by their estimates are presented in Figure E.3 for each residual series. They show that the marginal distributions of the series are reasonably well captured, although there is a single large outlier in the third residual series (which disappears if the March 2020 residual is removed; not shown). Therefore, the overall adequacy of the model seems, in our view, reasonable.
@@
Figure E.1: Auto- and crosscorrelation functions of the residuals of the fitted two-regime sixth-order LSTVAR model for the lags $0,1,...,24$. The lag zero autocorrelation coefficients are omitted, as they are one by convention. The dashed lines are the $95\%$ bounds ±$\pm 1.96/\sqrt{T}$ for autocorrelations of IID observations.
@@
Figure E.2: Auto- and crosscorrelation functions of the squared standardized residuals of the fitted two-regime sixth-order LSTVAR model for the lags $0,1,...,24$. The lag zero autocorrelation coefficients are omitted, as they are one by convention. The dashed lines are the $95\%$ bounds ±$\pm 1.96/\sqrt{T}$ for autocorrelations of IID observations.
@@
Figure E.3: Standardized residual time series and quantile-quantile-plots of the fitted two-regime sixth-order LSTVAR model. The quantile-quantile plots are based on $t$–distributions with zero mean, variance one, and degrees-of-freedom as well as skewness parameter values given by their estimates.